# Enabling Approximate Joint Sampling in Diffusion LMs

### v1

@@ B.1 EBM Experiments
EBM (Small) 1.33 1.85 1.89 5.26 5.76 6.18
EBM (Big) 1.61 2.08 2.08 5.27 5.74 6.18
Table 3: We report the runtime of the EBM algorithm (with Qwen-2.5-0.5B) for experimental setup in Table 1 in Tokens/sec. These numbers are comparable to those in figure 1. In this table we report the negative log-likelihood numbers on decoding with EBM. We fix the generation length to be 128 tokens and nucleus sampling to be 0.95, and temperature to be 1.0. Here we use Qwen2.5-7B as the verifier for EBM (Big), i.e., the same model which is used to evaluate responses, while using Qwen2.5-0.5B as the verifier for EBM (Small). We can see identical results because at each position we sample only two options both of which are identically rated by both the small and the big models
B.2 APD Experiments
B.2.1 NLL Experiments
We first argue that left-to-right decoding is not the right setting for generative perplexity. This can be seen as decoding multiple tokens in parallel for left-to-right decoding leads to worse negative log-likelihood (as measured by Qwen2.5-7B) than standard entropy based unmasking used by Dream. This is shown by the top table in Table 4. The bottom table shows when we use adaptive parallel decoding for unmasking. Using APD helps increase toks/step while decreasing NLL, but it still doesn’t outperform our joint sampler (or even marginal sampler) in NLL v/s Toks/Step plot (see Fig 4)
Spec=0 Spec=1 Spec=2 Spec=3
Marginal 4.39 7.94 8.28 8.53
R=0.3 R=0.5
NLL Tok/step Tok/sec NLL Tok/step Tok/sec
Marg + APD(U) 5.71 1.78 14.55 6.70 2.37 18.77
R=0.7 R=0.9
NLL Tok/step Tok/sec NLL Tok/step Tok/sec
Marg + APD(U) 7.74 4.13 31.31 8.51 10.54 76.92
Table 4: In this table we report the negative log-likelihood numbers on left-to-right generation. We fix the generation length to be 128 tokens and nucleus sampling to be 0.95, and temperature to be 1.0. Here we use Qwen2.5-0.5B as the verifier for APD.
B.2.2 GSM8K Experiments
We here recreate APD algorithm (Table 5) number for Dream-7B-Base model denoted by Mar+APD (U). The results are encapsulated in the caption.
R=0.3 R=0.5 R=0.7 R=0.9
Acc T/s Acc T/s Acc T/s Acc T/s
Base Mar + APD (U) 70.80 7.25 70.80 7.98 67.20 8.30 45.60 11.62
Mar + APD 71.20 3.29 72.40 3.39 71.20 3.44 65.20 3.65
Joint + APD 72.40 3.34 72.00 3.44 68.80 3.50 63.60 3.70
Table 5: In this table we adaptively decide the number of tokens to speculate by using the APD algorithm. $R$ denotes the APD hyper-parameter, where a larger value of $R$ puts more weight on the diffusion model (and hence more parallel generation) while a smaller value of $R$ puts more weight on AR model (and hence more sequential generation). T/s denotes tokens/step. We use Qwen2.5-0.5B to verify the answers. We compare the accuracy on the GSM8K dataset and the average number of tokens per steps for marginal sampling and joint sampling when combined with APD with different ratios $R$. Both, Mar+APD and Joint+APD denotes APD restricted to decoding a max of four tokens in each time step. Mar+APD(U), where U stands for unbounded denotes the standard APD algorithm.
B.3 Joint Sampler Loss
KL@1 KL@2 KL@3
Marginal 0.096 0.164 0.219
Joint 0.061 0.104 0.140
Table 6: In this table we report the KL loss for each speculation head, i.e., given base model features at time $t$ and spec additional ground truth tokens, we compute the KL divergence between the target logits and our predicted (for marginal, the base logits)
B.4 Architecture
@@
B.5 Grounding Probabilities
We also tried an addition by grounding the predicted probabilities in marginal probabilities. The basic idea can be stated as follows :
Suppose at some denoising step, the current input is $\bm{x}$} and the predicted marginals are $p(\cdot\mid{\bm{x}})$. In the main text we argued that after sampling a token (say at position $i$), the probability distribution of other positions in the sequence are not yet conditioned on the sampled output at position $i$. Hence sampling another token from $p(\cdot\mid{\bm{x}})$ would lead to an incorrect joint sample. But if the probability of the sampled token at position $i$ is close to $1$, i.e., $p^{i}(x^{i}\mid{\bm{x}})\approx 1$, then the marginal probabilities of other positions $p(\cdot\mid{\bm{x}})$ are already conditioned on this sample. A similar argument was made in  Wu et al. (2025) and proved using Boole’s inequality.
We can use this fact to ground our predicted probabilities in marginal probabilities. Suppose the probabilities after conditioning on token $x^{i}$ are $p^{*}$ and the original probabilities were $p_{0}$. Also let the probability of sampling $x^{i}$ be $c$, i.e. $p^{i}(x^{i}|{\bm{x}})=c$. Then we can say that
$c\cdot p^{*}+(1-c)\cdot\tilde{p}=p_{0}$
where $p^{*}$ is the probability $p(\cdot\mid{\bm{x}}\cup x^{i})$ and $p_{0}$ is the probability $p(\cdot\mid{\bm{x}})$, and $\tilde{p}$ is the probability $p(\cdot\mid{\bm{x}}\cup\{X^{i}\neq x^{i}\})$. Rearranging these terms gives us :
$p^{*}=\frac{p_{0}-\bigl((1-c)\cdot\tilde{p}\bigr)}{c}$
Now since $0\leq\tilde{p}\leq 1$, we get :
$\frac{p_{0}-(1-c)}{c}\leq p^{*}\leq\frac{p_{0}}{c}$
The same bound holds for our predictions of $p^{*}$. This can be implemented as a simple clamping operation. Note that the bound are tighter for $c\approx 1$ with $p^{*}=p_{0}$ for $c=1$.
Appendix C Algorithms
Algorithm 3 Marginal Sampling
1:Input: Prompt $\bm{x}$} , discrete diffusion model $f$ and the language model head $\bm{W}$}, tokens to unmask in each step $K$, unmasking logic TOP
2:while $|\mathcal{M}({\bm{x}})|>0$ do
3:  ${\bm{h}}\leftarrow f({\bm{x}})$ $\triangleright$ Forward pass through diffusion model
4:  for $k=1$ to $K$ do
5:   $i\leftarrow{\mathrm{TOP}}\{{\bm{W}}{\bm{h}}^{j}:j\in\mathcal{M}({\bm{x}})\}$ $\triangleright$ Select the position to be unmasked
6:   $\tilde{x}^{i}\sim\mathrm{softmax}({\bm{W}}{\bm{h}}^{i})$ $\triangleright$ Sample a token for that position
7:   ${\bm{x}}\leftarrow{\bm{x}}\oplus\tilde{x}^{i}$ $\triangleright$ Unmask token by repeating unmasking logic
8:  end for
9:end while
10:return $\bm{x}$}
@@ D.1 Sampling Parameters
Dream-7B-Base

### v2

@@ C.3 EDLM Experiments
EDLM (Small) 1.33 1.85 1.89 5.26 5.76 6.18
EDLM (Big) 1.61 2.08 2.08 5.27 5.74 6.18
Table 6: We report the runtime of the EDLM algorithm (with Qwen-2.5-0.5B) for experimental setup in Table 1 in Tokens/sec. These numbers are comparable to those in figure 1. In this table we report the negative log-likelihood numbers on decoding with EDLM. We fix the generation length to be 128 tokens and nucleus sampling to be 0.95, and temperature to be 1.0. Here we use Qwen2.5-7B as the verifier for EDLM (Big), i.e., the same model which is used to evaluate responses, while using Qwen2.5-0.5B as the verifier for EDLM (Small). We can see identical results because at each position we sample only two options both of which are identically rated by both the small and the big models
C.4 Architecture
@@ D.1 Sampling Parameters
Pretrained-only (Dream-Base/Llada-Base)
@@
Dream-7B-Coder
For HumanEval we use 0.2 temperature, 0.9 topp, and generation length of 512 and four few-shot examples. Similarly for MBPP we use the same sampling parameters
Appendix E Unmasking Logic
In our discussion till now, we assumed the list of positions to be generated $\sigma$ to be given to us. For the current state-of-the-art diffusion LLMs, positions to be generated next are decided by an operator which is a function of the logits, denoted as TOP. Base diffusion language models specify their criterion to decide the most suitable position to generate next. For example, Dream models (Ye et al., 2025) select the position with the least entropy to generate next, Llada (Nie et al., 2025) unmasks position with the highest confidence. ADJUST is agnostic to the choice of the unmasking logic and uses the base model’s unmasking logic for generation. A subtle difference to note is that naive parallel sampling of $K$ tokens calls TOP on the same output embeddings $f({\bm{x}})$, $K$ times, yielding top $K$ positions with (say) least entropy in ${\bm{W}}f$. ADJUST uses the updated logits to get the index of the next position, $\sigma(k)=\text{TOP}\left\{{\bm{W}}{\bm{h}}_{k}\right\}$, which we qualitatively observe are often different than using the base model’s logits ${\bm{W}}f$.